### v2

# Ondřej Mokrý Dept. of Telecommunications

Abstract
The paper focuses on inpainting missing parts of an audio signal spectrogram. The autoregression-based Janssen algorithm, the state-of-the-art for the time-domain audio inpainting, is adapted for the time-frequency setting. This novel method, termed Janssen-TF, is compared to the deep-prior neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
Index Terms:
audio inpainting, autoregression, deep prior, time-frequency, spectrogram.
I Introduction
Audio inpainting, sometimes referred to as missing audio interpolation or concealment, is the task of filling in the absent parts of audio with a meaningful content. Typically, the missing information is in the form of lost samples in the time domain. If compact regions of such samples are present, they are often referred to as the gaps. For the described task, plenty of methods have been developed: those based on the autoregressive character of audio [1, 2], methods based on the sparsity of the time-frequency representation of audio [3, 4, 5, 6], or methods relying on similarities in audio recordings [7, 8, 9]. Quite recently, deep learning audio inpainting methods appeared [10, 11]. For gaps of up to 50 ms in length, the Janssen method [2, 12] has been evaluated the best for a long time, but PHAIN [13] may become a new state-of-the art.
The time-domain inpainting represents the most natural case from the point of view of the application to signal recovery. Nonetheless, many practical algorithms, such as classic audio codecs (MP3, AAC, etc.), operate on short-time signal spectra. From this perspective, it is valuable to think about inpainting in the time-frequency domain. [^1] In other fields, one can go even further and require inpainting purely in the frequency domain, see for example [14], [15]. In such a case, a portion of time-frequency (TF) coefficients222 typically in the form of a complex-valued matrix are missing, which need to be recovered to allow a later reconstruction of the time-domain signal. Indeed, there have been a few methods developed in the TF context, all of them relying on neural modeling [16, 17, 18]. The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not need training in the most common sense.
Note that a compact gap in time bears some fundamental differences compared with a gap in the form of a missing number of consecutive columns of a spectrogram. This follows from the nature of time-frequency transforms, which use windows with overlaps to compute spectrograms, see for instance [19]. In particular, a gap in a spectrogram does not fully correspond to a time-domain gap and vice versa; there is always an amplitude fading in one of the domains thanks to the windowing.
The goal of this paper is to adapt the Janssen algorithm [2], proven successful in the time-domain case, for the TF-domain case, and compare its performance with the recent deep-prior-based approach of [18], both using objective metrics and by listening tests.
II Deep Prior Audio Inpainting (DPAI)
The deep prior, introduced in [20], is a concept used for the reconstruction of corrupted signals using a convolutional neural network (CNN). In contrast to most CNN applications, [20] does not utilize a training dataset; the reconstruction only involves a proper CNN architecture and the corrupted signal. The deep prior concept has also proven that the network architecture matters much more than previously thought.
II-A Application of Deep Prior to Audio Inpainting
The deep prior reconstruction can be utilized in a variety of applications, as outlined in [20]. In the context of the time-domain audio inpainting, consider a corrupted signal $\mathbf{x}^{\textup{cor}}\!\in\mathbb{R}^{N}$, having missing regions (filled with zeros), compared with the original signal $\mathbf{x}^{\textup{orig}}\!\in\mathbb{R}^{N}$. Formally, if a binary mask $\mathbf{m}\in\{0,1\}^{N}$ is utilized to represent the missing sections of the signal, the relation $\mathbf{x}^{\textup{cor}}=\mathbf{m}\odot\mathbf{x}^{\textup{orig}}$ holds, where $\odot$ represents the elementwise product. Such a relation naturally characterizes the feasible set $\mathit{\Gamma}_{\textup{T}}=\{\mathbf{x}\,|\,\mathbf{m}\odot\mathbf{x}=\mathbf{m}\odot\mathbf{x}^{\textup{orig}}\}$.
An encoder–decoder CNN denoted $f_{\theta}(\mathbf{n})$ is assumed to be able to generate any signal of size $N$ using a random but fixed noise input $\mathbf{n}\in\mathbb{R}^{N}$, using a proper $\theta$. This is a reliable assumption since the number of network parameters ${\theta}$ (weights and biases) is significantly greater than $N$.
In the deep prior reconstruction, parameters $\theta$ are initialized randomly. To obtain a reconstruction result, $\theta$ is then optimized as follows:
$\theta^{*}=\operatorname*{argmin}_{\theta}E(\mathbf{m}\odot f_{\theta}(\mathbf{n}),\mathbf{x}^{\textup{cor}}),$ (1)
where $E$ is a differentiable data fidelity function measuring the closeness of its two inputs. The optimal $\theta^{*}$ can be obtained using an iterative optimizer, e.g., Adam [21]. The crucial observation is that the optimization must be interrupted before it reaches $\theta^{*}$, for which we would have $f_{\theta^{*}\!}(\mathbf{n})=\mathbf{x}^{\textup{cor}}$, meaning basically overfitting. Instead, early stopping of the training is desired, resulting in an estimate $\hat{\theta}$ and a corresponding $\hat{\mathbf{x}}=f_{\hat{\theta}}(\mathbf{n})$ [20].
The authors of [18] used an analog approach to inpaint audio in the transformed domain. Through the use of the short-time Fourier transform (STFT), denoted $\mathbf{L}$ in the following, an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{L}\mathbf{x}=\mathbf{X}\in\mathbb{C}^{T\times F}$. Therefore the CNN operates on spectrograms and $\hat{\mathbf{X}}=f_{\hat{\theta}}(\mathbf{N})$ for a random but fixed matrix $\mathbf{N}$. Once $\hat{\mathbf{X}}$ has been established, the time-domain solution is obtained by the inverse STFT.
Regarding the mask, in [18] only the case of completely missing columns of the spectrogram is considered, i.e., for the TF mask $\mathbf{M}\in\{0,1\}^{T\times F}$ it holds that each column is either full of ones or zeros. Nevertheless, a generalization to other TF masks is straightforward. The observed spectrogram $\mathbf{X}^{\textup{cor}}$ is finally modeled as $\mathbf{X}^{\textup{cor}}\!=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}$, where $\mathbf{X}^{\textup{orig}}$ corresponds to $\mathbf{x}^{\textup{orig}}$ in the time domain, i.e., $\mathbf{X}^{\textup{orig}}=\mathbf{L}\mathbf{x}^{\textup{orig}}$. Note that by analogy with the time-domain case, we have the feasible set $\mathit{\Gamma}_{\textup{TF}}=\{\mathbf{X}\,|\,\mathbf{M}\odot\mathbf{X}=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}\}$.
The loss function $E$ in (1), which is optimized with respect to $\theta$, is in the form
$E=\alpha_{1}\mathrm{MSE}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}})+\alpha_{2}\mathrm{MSS}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}}),$ (2)
where MSE is the Mean Squared Error and $\mathrm{MSS}$ is the multi-scale spectrogram loss [22]. In [18], the weights $\alpha_{1}$ and $\alpha_{2}$ have values 1 and 0.1 respectively.
II-B Architecture
The U-Net is a well-known CNN architecture consisting of several convolutional layers, downsampling (encoder) and upsampling (decoder) operations. In certain applications including audio inpainting, the existence of skip connections between the encoder and decoder stages is important. The authors of [18] used an architecture similar to U-Net called MultiResUNet [23] with an additional inclusion of the so-called harmonic convolution [24], which proved to be the best among multiple tested architectures in [18]. The harmonic convolution is a variation on the common 2D convolution that takes into account higher harmonics of audio.
II-C Variants
The DPAI reconstructs the whole $\hat{\mathbf{X}}$. Yet, considering that only a part of the spectrogram was unavailable, the authors of [18] evaluate the performance of their method after injecting the reliable coefficients back to the solution as a postprocessing step. Formally, such an action is a projection
$\operatorname{proj}_{\mathit{\Gamma}_{\textup{TF}}}(\hat{\mathbf{X}})=\mathbf{M}\odot\mathbf{X}^{\textup{cor}}+(\mathbf{1}-\mathbf{M})\odot\hat{\mathbf{X}}$ (3)
onto the set $\mathit{\Gamma}_{\textup{TF}}$. Note that a resulting sharp transition between the regions is smeared in the time domain due to the nature of the STFT [19]. We call the reconstructions modified by (3) with context, while the others using the entire $\hat{\mathbf{X}}$ are without context. In the experimental part, we will study whether such a modification brings a significant performance improvement.
III Janssen spectrogram inpainting (Janssen-TF)
A signal $\mathbf{x}=[x_{1},\dots,x_{N}]^{\top}\!\in\mathbb{R}^{N}$ corresponds to an autoregressive process if the following holds:
$\sum_{i=1}^{p+1}a_{i}x_{n+1-i}=e_{n},\quad n=1,\dots,N+p,$ (4)
where $\mathbf{a}=[1,a_{2},\dots,a_{p+1}]^{\top}$ is the vector of model parameters (AR coefficients), $p$ is the model order, and $\mathbf{x}$ is zero-padded such that its final length is $N+p$. The error term $\mathbf{e}\in\mathbb{R}^{N+p}$ represents a realization of a zero-mean white noise process with variance $\sigma^{2}$ [25, Def. 3.1.2].
In the time-domain inpainting method proposed by Janssen et al. [2], the restored audio signal is supposed to follow the AR model (4) while both the missing samples in the time-domain and the unknown AR parameters $\mathbf{a}$ are being estimated. Janssen et al. approach the problem iteratively by performing two principal steps in the $i$-th iteration:
  1. 1.
    for a temporary solution $\mathbf{x}^{(i)}$, estimate the AR model parameters $\mathbf{a}^{(i)}$;
  2. 2.
    given the model parameters $\mathbf{a}^{(i)}$, estimate a new temporary solution $\mathbf{x}^{(i+1)}$ which fits the observed samples and the norm of the model error $\mathbf{e}$ is minimized.
Step 1 can be solved efficiently using standard methods, such as the Durbin–Levinson recursion (see e.g. the function lpc in Matlab). Regarding step 2, we propose to adapt the Janssen iteration to the TF setting, denoted Janssen-TF. In our adaptation, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, while the signal itself follows the estimated AR model. To formalize this modification, note that the error in (4) can be written as $\mathbf{e}=\mathbf{A}\mathbf{x}$, where $\mathbf{A}\in\mathbb{R}^{(N+p)\times N}$ is a Toeplitz matrix composed using the (fixed) coefficients $\mathbf{a}$. We can then pose the proposed modification of step 2 as
$\mathbf{x}^{(i+1)}=\arg\min_{\mathbf{x}}\tfrac{1}{2}\|\mathbf{A}\mathbf{x}\|^{2}+\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\mathbf{x}),$ (5)
where $\mathbf{L}$ is the STFT and $\iota_{\mathit{\Gamma}_{\textup{TF}}}$ denotes the indicator function of the convex set $\mathit{\Gamma}_{\textup{TF}}$ of feasible spectrograms. [^3] Note that the original Janssen algorithm minimizes the function $\frac{1}{2}\|\mathbf{A}\mathbf{x}\|+\iota_{\mathit{\Gamma}_{\textup{T}}}(\mathbf{x})$ in subproblem 2. As such, the problem (5) is convex and it is solvable using standard numerical methods. We propose to employ the alternating direction method of multipliers (ADMM), which includes two principal steps [26, Sec. 3.1.1]:
$\bar{\mathbf{x}}^{(k+1)}$ $=\arg\min_{\mathbf{x}}\tfrac{1}{2}\|\mathbf{A}\mathbf{x}\|^{2}+\tfrac{\rho}{2}\|\mathbf{L}\mathbf{x}-\mathbf{z}^{(k)}+\mathbf{u}^{(k)}\|^{2}\!,$ (6a)
$\mathbf{z}^{(k+1)}$ $=\arg\min_{\mathbf{z}}\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{z})+\tfrac{\rho}{2}\|\mathbf{L}\bar{\mathbf{x}}^{(k+1)}-\mathbf{z}+\mathbf{u}^{(k)}\|^{2}\!.$ (6b)
The whole algorithm is summarized in Alg. 1. For the simplification of (6a), we have posed the assumption [^4] This assumption is met for instance if $\mathbf{L}$ corresponds to a Parseval tight frame, which can be ensured by a suitable setting of the STFT. $\mathbf{L}^{*}\mathbf{L}=\mathbf{I}$, leading to $\bar{\mathbf{x}}^{(k+1)}=\operatorname{prox}_{\frac{1}{2}\|\cdot\|^{2}}(\mathbf{L}^{*}(\mathbf{z}^{(k)}-\mathbf{u}^{(k)}))$, see [27, Rem. 3], which then corresponds to line 1 of Alg. 1 [28, Table 1]. Similarly, (6b) boils down to line 1 of the algorithm; this is computed via (3). The hyperparameters of the algorithm are the number of outer and inner iterations ($I$ and $K$, respectively), and the step size $\rho>0$.
1 initialize $\mathbf{x}^{(1)}$
2 for $i=1,\dots,I$ do
       // AR model update
3       $\mathbf{a}^{(i)}=\texttt{lpc}(\mathbf{x}^{(i)})$
       // signal update using ADMM
4       form the Toeplitz matrix $\mathbf{A}$ from $\mathbf{a}^{(i)}$
5       $\mathbf{z}^{(1)}=\mathbf{L}\mathbf{x}^{(i)}$, $\mathbf{u}^{(1)}=\mathbf{0}$
6       for $k=1,\dots,K$ do
7             $\bar{\mathbf{x}}^{(k+1)}=(\mathbf{I}+\frac{1}{\rho}\mathbf{A}\!^{\top}\!\mathbf{A})^{-1}\mathbf{L}^{*}(\mathbf{z}^{(k)}-\mathbf{u}^{(k)})$
8             $\mathbf{z}^{(k+1)}=\operatorname{proj}_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\bar{\mathbf{x}}^{(k+1)}+\mathbf{u}^{(k)})$
9             $\mathbf{u}^{(k+1)}=\mathbf{u}^{(k)}+\mathbf{L}\bar{\mathbf{x}}^{(k+1)}-\mathbf{z}^{(k+1)}$
10       end for
11      $\mathbf{x}^{(i+1)}=\bar{\mathbf{x}}^{(K+1)}$
12
13 end for
Algorithm 1 Janssen-TF-ADMM
Note that several algorithms can be used to solve (5) numerically, such as the primal–dual algorithm [29] or the Douglas–Rachford or the forward–backward algorithms [28], combined with the approximal operator of the function $\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\,\cdot)$ [30]. However, the variant with ADMM described above, denoted Janssen-TF-ADMM, exhibited the best performance out of six algorithms in our preliminary testing.
IV Experiments
IV-A Data, setup and metrics
For our first experiment, we adopt the audio dataset [^5] https://github.com/fmiotello/dpai/ used for the evaluation of DPAI [18]. It consists of 6 music and 2 speech recordings of 5 seconds in length, sampled at 16 kHz. Only minor modifications were made to the original distribution of gaps (i.e., the placement of the missing spectrogram columns), such that each test excerpt includes 5 gaps of the same length. The gap lengths range from 1 to 6 missing columns. In turn, the entire test set comprises $8\times 6=48$ corrupt spectrograms. Even though the dataset is relatively small and includes only a few types of audio signals, it provides a fair comparison of the proposed method with DPAI—the choice allows the use of exactly the same U-net architecture of DPAI from [18]. In particular, we chose their “best2” setting, and 5000 epochs of the Adam optimizer.
To validate the results on a larger dataset, we hand-picked 60 music recordings of various complexity from the IRMAS dataset [31, 32]. To enable the use of DPAI, the recordings were subsampled to 16 kHz and shortened to 5 seconds.
The STFT (i.e., the operator $\mathbf{L}$) utilizes a 2048-sample-long Hann window (which corresponds to 128 ms), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^6] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
We compare the proposed method Janssen-TF-ADMM with three baseline methods:
  • DPAI reconstruction with context,
  • DPAI reconstruction without context,
  • gapwise Janssen method in the time domain [12].
The gapwise method is a twist to the original Janssen algorithm [2], which, instead of processing by means of overlapping frames, makes use of a single segment consisting of the whole gap and its nearby context.
To objectively evaluate the reconstruction quality, we assess the difference between the inpainted signal $\hat{\mathbf{x}}$ and the clean reference $\mathbf{x}^{\textup{orig}}$ using two standard metrics:
  1. 1.
    signal-to-noise ratio (SNR), defined in decibels as $\text{SNR}(\hat{\mathbf{x}},\mathbf{x}^{\textup{orig}})=10\log_{10}\frac{\|\mathbf{x}^{\textup{orig}}\|^{2}}{\|\mathbf{x}^{\textup{orig}}-\hat{\mathbf{x}}\|^{2}}$ [3],
  2. 2.
    objective difference grade (ODG) computed using the PEMO-Q [33], which takes psychoacoustics into account and predicts the subjective evaluation of the difference between $\hat{\mathbf{x}}$ and $\mathbf{x}^{\textup{orig}}$ on the scale from $-4$ to 00 (from “very annoying” to “imperceptible”).
In addition, we run
  1. 3.
    a subjective listening test containing a limited number of excerpts. The MUSHRA-type test [34] was performed, which includes the reference signal, the competing recosntructions, and finally the hidden anchor, i.e., the signal reconstructed from the spectrogram with gaps. The participants were asked to score the perceived similarity of the signals with the reference signal on a continuous scale from 0 to 100, with the help of the webMUSHRA environment [^7] https://github.com/audiolabs/webMUSHRA/. The test was run in a quiet music studio, using a professional sound card and headphones. The listening conditions were identical for all the participants. Following the recommendation [34], 11 assessors out of 18 were selected through post-screening.
IV-B Results
The objective comparison, including the proposed Janssen-TF method, is shown in Fig. 1 for the case of the DPAI dataset. The first observation is that the time-domain-based Janssen method [12] is not competitive with the other approaches. This can be explained by the incompatibility of gaps in the time and time-frequency domains (a sharp mask in the spectrogram domain affects a longer signal portion when transformed into the time domain, and the whole affected segment is treated as missing). While the average performance of the DPAI variants mostly lies within the confidence intervals of each other, suggesting limited significance of the differences, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except the longest gaps).
Figure 1: Comparison of the inpainting methods using objective metrics – SNR (top) and ODG (bottom). The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [35, p. 160]. If the intervals do not overlap, it may be concluded that the difference of the means is statistically significant.
To support the objective evaluation on the first dataset, Fig. 2 presents the results of the subjective listening test. Its results correlate well with the objective evaluation; in particular, Janssen-TF-ADMM is superior to the other methods. Moreover, this result is statistically significant in terms of medians. On the other hand, the DPAI variants scored almost the same.
anchorreferenceDPAI without contextDPAI with contextgapwise JanssenJanssen--TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score
Figure 2: A boxplot showing the distribution of scores in the listening test. The individual boxes span from the 25th to the 75th percentile of the recorded scores. The notches (filled areas) around the medians (orange lines) are constructed such that boxes whose notches do not overlap have different medians at the 5% statistical significance level.
Finally, Fig. 3 shows the evaluation on the IRMAS subset comprising 60 signals. This evaluation supports the previous analysis and leads to the conclusion that the proposed method is superior to the others. For shorter gaps, DPAI with context is preferable over DPAI not exploiting context.
Figure 3: SNR and ODG results on the IRMAS dataset.
Regarding computational demands, a single run of DPAI takes ca 19 minutes on the NVIDIA Tesla V100S GPU with 32 GB of memory, regardless of the number of missing spectrogram columns. Reconstructing a single signal with Janssen-TF-ADMM takes 10 to 20 minutes, proportional to the number of columns; this performance corresponds to a PC with Intel Core i7 3.40 GHz processor, 32 GB RAM.
V Conclusion and Outlook
The paper has shown that the proposed method of spectrogram inpainting, Janssen-TF, performs significantly better than the recently introduced DPAI algorithm, which is based on the deep prior idea. The conclusion has been certified both by objective and by subjective tests.
Matlab source codes for Janssen-TF and other supplementary material are publicly available at https://github.com/rajmic/spectrogram-inpainting.